Evaluate the circuit shown at the right. What is the charge of capacitor C3? 20V C1=6 C2=6 C4= 2 C6=1


1
Expert's answer
2021-06-08T09:35:12-0400

"C_{12} = C_1 + C_2 = 12\u00b5F"

"C_{eq} =[ \frac{1}{C_{12}} + \frac{1}{C_{13}}]^{−1} = 1.5µF."

"Q_2 = V_2C_2 = 8\u00b5C"